Question text

Suppose the government imposes a tax of 20 percent on the first $40,000 of income and a tax of 30 percent on income above $40,000 but less than $90,000. Income over $90,000 is taxed at a rate of 40 percent. The marginal tax rate for a person whose income is $45,650 is 30 percent.

Select one:

True

False

Question text

Alvin earned $30,000 and paid $6,000 in taxes. If Beatrice earned $25,000 then she would have to pay less than $5,000 in taxes if the tax was progressive.

Select one:

True

False

Question text

A tax system in which higher-income taxpayers have the same tax rates, even though they pay a larger amount of tax, than lower-income taxpayers is a progressive tax system

Select one:

True

False
Answer #1

Answer
True
A marginal tax rate is the tax on the extra $1 earned
if the income increases from $45650 to $45651 then the tax on the extra dollar is 30% as it is in the bracket of a tax of 30 percent on income above $40,000 but less than $90,000.
=======
Answer
Tax is progressive if the average tax rate increases as the income increase
average tax rate =(tax amount /income)*100
ATR=(6000/30000)*100=20%
ATR<(5000/25000)*100=20%

the tax is progressive as the ATR is decreasing as income decreases.
==========
Answer
False
The amount is not considered, it is the average tax rate considered which is constant over so it is a flat tax system.